# Relativistic Leaky Surface Plasmon Accelerator in Nanostructured Carbon Nanotube Forest

### v1

@@ IV.3 Wall electron transition and cascaded leaky RSPs excitation
The key difference between the solid-state forest target and a uniform gaseous plasma is the presence of the solid inner surface, over which the electrical potential restricts the wall electrons from freely crossing the boundary. For an electron at the wall to overcome this potential barrier, it must have sufficient transverse momentum. Wall electrons gain greater momentum in stronger laser fields, increasing the probability of crossing the inner surface and becoming trapped in the leaky field. Therefore, for a specific target configuration, the parameter $a_{0}$ determines both the excitation of RSPs and the probability of wall electrons transitioning across the surface for injection.
@@
A series of simulations were performed with $a_{0}$ values ranging from $0.1$ to $220$. The effects on the amplitude of the acceleration field ($E_{z}$), the total charge of the injected electron beam ($Q_{e}$), and the mean energy ($\mathcal{E}$) of the accelerated electron beam are presented in Fig.8 (a), (b) abd (c), respectively. As $a_{0}$ increases, the RSPs become stronger, resulting in electron beams with higher charge that can be accelerated to greater energies. Edge injection occurs for $94<a_{0}<$ 9, while self-injection begins at $a_{0}\geq 9$, where a significantly larger number of electrons are injected into the leaky field. At $a_{0}>30$, the beam loading effect becomes significant, reducing the effective acceleration field and energy gain rate. At low values of $a_{0}<4$, wall electrons cannot overcome the inner surface potential barrier, preventing any electron injection into the central vacuum channel. When $a_{0}<2$, SP becomes non-relativistic. Simulations also reveal critical thresholds for bubble wakefield excitation and the blowout phenomenon. In a moderately high laser pulse range ($5030<a_{0}<$ 50), all the wall electrons gain sufficient transverse momentum to freely cross the inner surface. These results in the formation of plasma bubbles, similar to those in uniform plasmas, but with significantly enhanced field amplitudes that exceed the coherent field limit due to the mitigation of wave breaking. As the laser strength increases further ($a_{0}>50$), wall electrons begin to acquire enough outward momentum from ponderomotive scattering to cross the outer surface of the forest target of finite wall thickness and become ejected from the target—a phenomenon referred to as the blowout effect here. In the range $20050<a_{0}<$ 200, partial blowout occurs where a fraction of the wall electrons are ejected from the outer wall, leading to a reduction in the acceleration field ($E_{z}$), total injected charge ($Q_{e}$), and beam energy ($\mathcal{E}$). Full blowout occurs at extremely high laser strengths ($a_{0}>200$), where all wall electrons are permanently ejected from the outer surface. Consequently, no sustained acceleration field is formed within the vacuum channel, and particle acceleration ceases entirely. These thresholds for $a_{0}$ depend on the balance between laser pulse intensity and target geometry, varying with different target and laser configurations. This interplay highlights the importance of precise tuning to optimise the acceleration process while avoiding undesirable blowout effects.
@@
The robustness of RSP excitation enables efficient RSP-based acceleration in a cascading manner. This approach is highly advantageous for high-power laser-driven experiments, where both beam charge and energy can be significantly increased. Additionally, the energy spectrum of the accelerated electrons can be substantially improved. A high-intensity laser pulse can initially drive strong edge injection and self-injection. As the laser energy depletes and the laser strength ($a_{0}$) decreases, the self-injection process gradually ceases. However, the leaky field remains well-structured and stable, ensuring that the self-trapped electrons continue to be accelerated. This leads to a narrower energy spectrum for the accelerated beam. For example, a laser pulse with an initial strength of $a_{0}=30$ can trap an electron beam of charge $2.15\leavevmode\nobreak\ $\mathrm{n}\mathrm{C}$$ in the leaky field, which is accelerated to $180\leavevmode\nobreak\ $\mathrm{M}\mathrm{e}\mathrm{V}$$ after $354\leavevmode\nobreak\ $\mu\mathrm{m}$$ propagation as shown in Fig. 9. The energy spectrum has a $20\%$ spread and can extend up to $340\leavevmode\nobreak\ $\mathrm{M}\mathrm{e}\mathrm{V}$$. This indicates an average acceleration gradient of $G_{\text{AVG}}=0.53\leavevmode\nobreak\ $\mathrm{T}\mathrm{e}\mathrm{V}\mathrm{/}\mathrm{m}$$ and a peak of $G_{\text{PEAK}}=0.96\leavevmode\nobreak\ $\mathrm{T}\mathrm{e}\mathrm{V}\mathrm{/}\mathrm{m}$$.
@@
IV.4 Target geometry and simulation method
Table 2: Parameters of the porous CNT target used in PIC simulations.
Layer gap, $g$ $50\leavevmode\nobreak\ $\mathrm{n}\mathrm{m}$$
Pore diameter, $d$ $100\leavevmode\nobreak\ $\mathrm{n}\mathrm{m}$$
Forest channel radius, $r_{t}$ $3\leavevmode\nobreak\ $\mu\mathrm{m}$$
Filling factor, $f_{f}$ $0.8$
Number of layers, $N$ $20$
[…]
Vacuum sections of $20\leavevmode\nobreak\ $\mu\mathrm{m}$$ and $10\leavevmode\nobreak\ $\mu\mathrm{m}$$ are placed longitudinally at the head and tail of the forest target, respectively, for laser initialisation and electron beam diagnostics. The target species consists of carbon atoms, which can only be singly ionised to produce $C^{+}$ ions. Field ionisation is implemented using the Ammosov-Delone-Krainov (ADK) method  [66]. The dimensions of the moving window are $12\leavevmode\nobreak\ $\mu\mathrm{m}$\times 12\leavevmode\nobreak\ $\mu\mathrm{m}$\times 20\leavevmode\nobreak\ $\mu\mathrm{m}$$, comprising $384\times 384\times 512$ cells in the $x$, $y$, and $z$ directions, respectively. This configuration provides 3 points within each forest pore transversely, ensuring an accurate spatial representation of the structure. Each cell contains 8 macro particles, which is sufficient to solve the electron dynamics effectively. For particle and field updates, the simulations utilise the Boris pusher [67] and the CKC Maxwell solver [68].
V Conclusion and summary
@@
In this paper, we demonstrated that a TW optical laser pulse can effectively drive the excitation of strong RSPs on the $\mu m$-scale cylindrical surface of a CNT forest. These RSPs feature an electromagnetic (EM) component that leaks into the central vacuum channel, providing acceleration fields on the order of $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$. These fields are suitable for accelerating both relativistic negatively and positively charged particles. We propose two distinct electron injection schemes, each with unique features that enhance the experimental feasibility of this approach, paving the way for practical implementations. We also analyzed the fundamental physical principles that enable this mechanism to function. It requires the CNT forest to be in a quantum-degenerate and correlated state, a condition that is achievable under standard experimental setups. This provides a solid theoretical foundation for the feasibility of the proposed approach.
@@
In line with recent advances in both laser-driven solid-state and plasma-based accelerators, this paper presents a novel concept for accelerating both positively and negatively charged particles by exciting a new class of SP modes, RSPs, through the ponderomotive scattering of a high-intensity optical laser pulse passing paraxially through a nanostructured CNT forest. Compared to conventional conductive materials, such as metals and semiconductors, CNTs can offer significantly higher plasmon gain due to their unique electronic structure and exceptional field confinement. This enables stronger light-matter interactions, which can substantially improve energy efficiency in experimental setups. Unlike traditional solid-state accelerators, such as crystal-based accelerators or dielectric laser accelerators (DLAs), the CNT-based accelerator leverages plasma dynamics, promising exceptionally high acceleration gradients of up to hundreds of $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$ in principle. This approach can generate lepton beams in the $\mathrm{M}\mathrm{e}\mathrm{V}$-level on $\mu\mathrm{m}$ scales or $\mathrm{P}\mathrm{e}\mathrm{V}$-level on meter scales, offering immense potential for future high-energy applications. In contrast to gaseous plasma-based accelerators, the CNT-based method is solid-state, capable of providing much higher acceleration fields with enhanced mechanical stability, making it a promising alternative for next-generation acceleration technologies.
@@
The experiments can be designed with the laser facilities of $\mathrm{T}\mathrm{W}$ to $\mathrm{P}\mathrm{W}$ power with $\mu\mathrm{m}$-scale beam size [69]. Recent advances in modern nanofabrication techniques have enabled the production of nanostructured CNT forests with specifications that can meet the experimental requirements [70]. This unique combination makes CNT-based particle accelerators a promising solution for pushing the energy frontier in scientific discovery and also poised to significantly enhance our precision in manipulating matter for potential applications in cutting-edge fields like high-energy plasmonics, ion channeling and radiation generation.
acknowledgements
@@
The datasets used and/or analysed during the current study available from the corresponding author on reasonable request.
Author contributions
All authors conceived the numerical simulations and discussed the results. B.F.L. run the simulations, processed the data and wrote the manuscript. C.W. provided the computational resources. All authors reviewed and edited the manuscript.
Additional Information
The authors declare no competing interests.

### v2

@@ III.2 Self injection
Here, it is seen that the key difference between the structured CNT plasma and a uniform gaseous plasma is the presence of the solid inner surface, over which the electrical potential restricts the wall electrons from freely crossing the boundary. For an electron on the wall to overcome this potential barrier, it must have sufficient transverse momentum. Wall electrons gain greater momentum in stronger laser fields, increasing the probability of crossing the inner surface and becoming trapped in the leaky field. Therefore, for a specific target configuration, the parameter $a_{0}$ determines both the excitation of SPs and the probability of wall electrons transitioning across the surface for injection.
@@
A series of simulations were performed with $a_{0}$ values ranging from $0.1$ to $220$. The effects on the amplitude of the acceleration field ($E_{z}$), the total charge of the injected electron beam ($Q_{e}$), and the mean energy ($\mathcal{E}$) of the accelerated electron beam are presented in Fig.7 (a)-(c), respectively. As $a_{0}$ increases, the SPs become stronger, resulting in electron beams with higher charge that can be accelerated to greater energies. Edge injection occurs for $94<a_{0}<$ 9, while self-injection begins at $a_{0}\geq 9$, where a significantly larger number of electrons are injected into the leaky field. At $a_{0}>30$, the beam loading effect becomes significant, reducing the effective acceleration field and energy gain rate. At low values of $a_{0}<4$, wall electrons cannot overcome the inner surface potential barrier, preventing any electron injection into the central vacuum channel. When $a_{0}<2$, SP becomes non-relativistic. Simulations also reveal critical thresholds for bubble wakefield excitation and the blowout phenomenon. In a moderately high laser pulse range ($5030<a_{0}<$ 50), all the wall electrons gain sufficient transverse momentum to freely cross the inner surface. This results in the formation of plasma bubbles, similar to those in uniform plasmas, but with significantly enhanced field amplitudes that exceed the coherent field limit due to the mitigation of wave breaking. As the laser strength increases further ($a_{0}>50$), wall electrons begin to acquire enough outward momentum from ponderomotive scattering to cross the outer surface of the forest target of finite wall thickness and become ejected from the target—a phenomenon referred to as the blowout effect here. In the range $20050<a_{0}<$ 200, partial blowout occurs where a fraction of the wall electrons are ejected from the outer wall, leading to a reduction in the acceleration field ($E_{z}$), total injected charge ($Q_{e}$), and beam energy ($\mathcal{E}$). Full blowout occurs at extremely high laser strengths ($a_{0}>200$), where all wall electrons are permanently ejected from the outer surface. Consequently, no sustained acceleration field is formed within the vacuum channel, and particle acceleration ceases entirely. These thresholds for $a_{0}$ depend on the balance between laser pulse intensity and target geometry, varying with different target and laser configurations. This interplay highlights the importance of precise tuning to optimise the acceleration process while avoiding undesirable blowout effects.
@@
The robustness of SP excitation enables efficient SP-based acceleration in a cascading manner. This approach is highly advantageous for high-power laser-driven experiments, where both beam charge and energy can be significantly increased. Additionally, the energy spectrum of the accelerated electrons can be substantially improved. A high-intensity laser pulse can initially drive strong edge injection and self-injection. As the laser energy depletes and the laser strength ($a_{0}$) decreases, the self-injection process gradually ceases. However, the leaky field remains well-structured and stable, ensuring that the self-trapped electrons continue to be accelerated. This leads to a narrower energy spectrum for the accelerated beam. For example, a laser pulse with an initial strength of $a_{0}=30$ can trap an electron beam of charge $2.15\leavevmode\nobreak\ $\mathrm{n}\mathrm{C}$$ in the leaky field, which is accelerated to $180\leavevmode\nobreak\ $\mathrm{M}\mathrm{e}\mathrm{V}$$ after $354\leavevmode\nobreak\ $\mu\mathrm{m}$$ propagation as shown in Fig. 8. The energy spectrum has a $20\%$ spread and can extend up to $340\leavevmode\nobreak\ $\mathrm{M}\mathrm{e}\mathrm{V}$$. This indicates an average acceleration gradient of $G_{\text{AVG}}\approx 0.51\leavevmode\nobreak\ $\mathrm{T}\mathrm{e}\mathrm{V}\mathrm{/}\mathrm{m}$$ and a peak of $G_{\text{PEAK}}\approx 0.96\leavevmode\nobreak\ $\mathrm{T}\mathrm{e}\mathrm{V}\mathrm{/}\mathrm{m}$$.
@@
IV Conclusion and summary
@@
In this paper, we propose that a TW optical laser pulse can effectively drive the excitation of strong SPs on the $\mu m$-scale cylindrical surface nanostructured inside a CNT forest. These SPs feature an EM component that leaks into the central vacuum channel, providing acceleration fields on the order of sub-$\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$. These fields are suitable for accelerating both relativistic negatively and positively charged particles. We present two distinct electron injection schemes, each with unique features that enhance the experimental feasibility of this approach, paving the way for practical implementations.
@@
In line with recent advances in both laser-driven solid-state and plasma-based accelerators, this paper presents a novel concept for accelerating both positively and negatively charged particles by exciting non-evanescent leaky EM fields of SP modes through the ponderomotive scattering of a high-intensity optical laser pulse passing paraxially through a nanostructured CNT forest. Compared to conventional conductive materials, such as metals and semiconductors, CNTs can offer significantly higher plasmon gain due to their unique electronic structure and exceptional field confinement. This enables stronger light-matter interactions, which can substantially improve energy efficiency in experimental setups. Unlike traditional solid-state accelerators, such as crystal-based accelerators or dielectric laser accelerators (DLAs), the CNT-based accelerator leverages plasma dynamics, promising exceptionally high acceleration gradients of up to hundreds of $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$ in principle and avoinding the limitation from the channel size of natural crystals. This approach can generate lepton beams in the $\mathrm{M}\mathrm{e}\mathrm{V}$-level on $\mu\mathrm{m}$ scales or $\mathrm{P}\mathrm{e}\mathrm{V}$-level on meter scales, offering immense potential for future high-energy applications. In contrast to gaseous plasma-based accelerators, the CNT-based method is solid-state, capable of providing much higher acceleration fields with enhanced mechanical stability, making it a promising alternative for next-generation acceleration technologies.
@@
The experiments can be designed with the laser facilities of $\mathrm{T}\mathrm{W}$ to $\mathrm{P}\mathrm{W}$ power with $\mu\mathrm{m}$-scale beam size [38]. Recent advances in modern nanofabrication techniques have enabled the production of structured CNT forests with specifications that can meet the experimental requirements [39]. This unique combination makes CNT-based particle accelerators a promising solution for pushing the energy frontier in scientific discovery and also poised to significantly enhance our precision in manipulating matter for potential applications in cutting-edge fields like high-energy plasmonics, ion channelling and radiation generation.
Acknowledgements.
@@
The datasets used and/or analysed during the current study are available from the corresponding author on reasonable request.